. At what distance along the central axis of a ring of radius R and uniform charge Q is the magnitude of the electric field maximum?


1
Expert's answer
2021-07-04T17:40:45-0400

We know that

"E=\frac{kqx}{(x^2+R^2)^\frac{3}{2}}"

"\frac{dE}{dy}=0"

"\frac{dE}{dy}=Kq\frac{(x^2+R^2)^\frac{3}{2}-\frac{3}{2}(x^2+R^2)^\frac{1}{2}.2x^2}{x^2+R^2}"

"Kq\frac{(x^2+R^2)^\frac{3}{2}-\frac{3}{2}(x^2+R^2)^\frac{1}{2}.2x^2}{x^2+R^2}=0"


"{(x^2+R^2)^\frac{3}{2}-\frac{3}{2}(x^2+R^2)^\frac{1}{2}.2x^2}=0"

"x^2+R^2=3x^2"

"2x^2=R^2"

"x=\frac{R}{\sqrt{2}}"

When distance "x=\frac{R}{\sqrt{2}}" Electric field is maximum